# MobileCLIP: Fast Image-Text Models through Multi-Modal Reinforced Training

### v1

@@ 5.3 Learning Efficiency
80.1 69.6 74.5 92.3 46.7 66.5
67.3
TinyCLIP-RN19M [67] LAION-400M [50]
15.2B
ResNet-19M
Custom
18.6 + 44.8
1.9 + 1.9
56.3 43.6 58.0 75.4 30.9 47.8
48.3
TinyCLIP-RN30M [67] LAION-400M [50]
15.2B
ResNet-30M
Custom
29.6 + 54.2
2.6 + 2.6
59.1 45.7 61.5 80.1 33.8 51.6
50.2
TinyCLIP-40M/32 [67] LAION-400M [50]
15.2B
ViT-40M/32
Custom
39.7 + 44.5
3.0 + 1.9
59.8 46.5 59.1 76.1 33.5 48.7
51.2
MobileCLIP-S0 DataCompDR-1B
13B
MCi0
MCt
11.4 + 42.4
1.5 + 1.6
67.8 55.1 67.7 85.9 40.4 58.7 58.1
OpenAI-RN50 OpenAI-400M [46]
13B
ResNet-50
Base
38.3 + 63.4
3.3 + 3.3
59.8 45.1 57.4 80.0 28.5 48.8
48.1
TinyCLIP-61M/32 [67] LAION-400M [50]
15.2B
ViT-61M/32
Custom
61.4 + 54.0
4.3 + 2.6
62.4 48.7 62.6 78.7 36.5 52.8
53.0
@@
15.8B
ViT-63M/32
Custom
64.5 (-) 66.0 84.9 38.5 56.9
MobileCLIP-S1 DataCompDR-1B
13B
MCi1
Base
21.5 + 63.4
2.5 + 3.3
72.6 60.7 71.0 89.2 44.0 62.2 61.3
OpenAI-RN101 OpenAI-400M [46]
13B
ResNet-101
Base
56.3 + 63.4
4.3 + 3.3
62.3 48.5 58.0 79.0 30.7 49.8
50.3
OpenAI-B/32 OpenAI-400M [46]
13B
ViT-B/32
Base
86.2 + 63.4
5.9 + 3.3
63.3 48.5 58.8 78.9 30.4 50.1
52.5
LAION-B/32 LAION-2B [51]
32B
65.7 51.9 66.4 84.4 39.1 56.2
54.8
DataComp-B/32 DataComp-1B [18]
13B
69.2 55.2 61.1 79.0 37.1 53.5
58.0
DataComp-B/32-256 DataComp-1B [18]
34B
ViT-B/32-256
Base
86.2 + 63.4
6.2 + 3.3
72.8 58.7 64.9 84.8 39.9 57.9
60.9
MobileCLIP-S2 DataCompDR-1B
13B
MCi2
Base
35.7 + 63.4
3.6 + 3.3
74.4 63.1 73.4 90.3 45.4 63.4 63.7
VeCLIP-B/16 [31] WIT-200M
6.4B
ViT-B/16
Base
86.2 + 63.4
11.5 + 3.3
64.6 (-) 76.3 91.1 48.4 67.2
OpenAI-B/16 WIT-400M [46]
13B
Base
86.2 + 63.4
11.5 + 3.3
68.3 55.9 67.7 85.9 40.4 58.7
58.1
LAION-B/16 LAION-2B [51]
34B
Base
86.2 + 63.4
11.5 + 3.3
70.2 56.6 69.8 86.3 42.3 59.4
58.7
EVA02-B/16 Merged-2B [54]
8B
Base
86.2 + 63.4
74.7 59.6 71.5 86.0 42.2 58.7
58.9
DFN-B/16 DFN-2B [16]
13B
Base
86.2 + 63.4
11.5 + 3.3
76.2 62.3 69.1 85.4 43.4 60.4
60.9
DataComp-B/16 DataComp-1B [18]
13B
Base
86.2 + 63.4
11.5 + 3.3
73.5 60.8 69.8 86.3 42.3 59.4
61.5
SigLIP-B/16 [76] Webli-1B
40B
Custom
92.9 + 110.3
9.9 + 5.8
76.0 61.0 74.7 89.1 47.8 65.7
62.3
MobileCLIP-B DataCompDR-1B
13B
Base
86.3 + 63.4
10.4 + 3.3
76.8 65.6 77.3 91.4 50.6 68.8 65.2
Table 6: MobileCLIP family of models has the best average performance at various latencies. Retrieval performances are reported @1. Last column shows average performance on 38 datasets as in OpenCLIP [28]. Models are grouped by their total latency in increasing order and by performance within each group. “Base” refers to standard CLIP Transformer-based [62] text encoder with 12 layers, and “Custom” stands for customized text encoder used in the respective method. For TinyCLIP-63M/32 and EVA02-B/16, we were unable to reliably benchmark models. Note: EVA02-B/16 [54] uses MIM pretrained weights for its vision encoder and OpenCLIP-B pretrained weights for its text encoder. TinyCLIP models use advanced weight initialization methods utilizing OpenCLIP models trained on LAION-2B[51] dataset. All other models, including ours are trained from scratch.
@@ 5.5 Retrieval Performance Analysis
Table 7: Performance on ARO benchmark. All the models use ViT-B/16 as image encoder and the standard CLIP text encoder. For VG Rel. and VG Attr. datasets, Macro Accuracy is reported and for Flickr30k-Order and COCO-Order recall@1 is reported. For more details please refer to [74].
@@ 6 Conclusion
Appendix A Experimental Setup
Additional details of our training and evaluation are provided here. We train all models at 224 resolution.
Table 10 summarizes the hyperparameters we used to train MobileCLIP-B on DataCompDR-1B. For other variants of MobileCLIP (S0, S1, and S2) we use the same hyperparameters except using $\lambda=1.0$. For experiments on DataCompDR-12M we use global batch size of 8192.
For our ensemble distillation ablations in Appendix C, we use 32 total A100 GPUs but we use the same global batch size of 8192 as our other ablations. We also use a smaller uniformly sampled DataComp-8M for ablations in Appendices C and B that results in a slightly lower performance than DataCompDR-12M used for the rest of ablations.
For ImageNet-1k experiments in Sec. 4.2, we follow the training recipe prescribed in [58, 37], i.e. the models are trained for 300 epochs using AdamW optimizer with weight decay of 0.05 and peak learning rate $10^{-3}$ for a total batch size of 1024. The number of warmup epochs is set to 5 and cosine schedule is used to decay the learning rate. The teacher model for distillation is RegNetY-16GF [47] Our implementation uses Timm library [65] and all the models were trained on single machine with 8$\times$×NVIDIA A100 GPUs. The hyperparameters for the three variants of MCi are detailed in Tab. 8. The performance of MCi variants is detailed in Tab. 9 and compared against recent state-of-art efficient architectures. MCi obtains the best trade-off amongst recent efficient architectures as seen in Fig. 7.
@@ Appendix A Experimental Setup
Appendix B Image Augmentation
@@
Appendix C CLIP Ensembles
@@
Appendix D Ablations on Lossy Compressions
In Tab. 2(c) we presented the storage sizes for DataCompDR-12M and DataCompDR-1B with BFloat16 compression of the embeddings. In this section, we further analyze the storage reduction by i) reducing the number of augmentations, and ii) lossy compression of embeddings.
@@ Appendix D Ablations on Lossy Compressions
Image Text Syn.
Aug. Params
Text Emb.
Image Emb.
BFloat16
Sparsity
Size (GBs)
@@
Appendix E Hybrid Text Encoder
@@
Appendix F Extended Results
@@
MobileCLIP-B 76.8 58.7 89.6 41.4 64.5 69.8 69.4 91.4 99.1 99.9 77.3 94.4 96.7 68.8 88.3 92.9 50.6 74.9 82.9
MobileCLIP-S2 74.4 49.3 87.0 46.9 62.2 66.8 66.6 90.3 98.9 99.6 73.4 92.3 95.6 63.4 85.1 91.4 45.4 70.1 79.0
MobileCLIP-S1 72.6 40.3 84.7 50.5 60.3 64.9 63.4 89.2 98.0 99.5 71.0 91.3 95.3 62.2 84.3 90.1 44.0 68.9 77.7
MobileCLIP-S0 67.8 26.5 78.6 53.8 55.5 59.9 55.9 85.9 97.1 98.8 67.7 88.8 93.3 58.7 81.1 88.2 40.4 66.0 75.9
@@
VeCLIP-B/16 [31] 64.6 (-) (-) (-) (-) 57.7 (-) 91.1 98.5 99.7 76.3 93.5 96.4 67.2 87.3 92.7 48.4 73.3 81.8
TinyCLIP-63M/32 [67] 64.5 22.8 74.1 (-) 50.8 55.7 31.2 84.9 (-) (-) 66.0 (-) (-) 56.9 (-) (-) 38.5 (-) (-)
CLIPA-B/16 [33] 63.2 26.8 73.2 (-) 48.7 55.6 44.3 75.3 (-) (-) 58.3 (-) (-) 53.1 (-) (-) 35.2 (-) (-)

### v2

@@ 5.3 Learning Efficiency
(-) (-) 80.1 69.6 74.5 92.3 46.7 66.5 67.3
TinyCLIP-RN19M [68] LAION-400M [51] 15.2B ResNet-19M Custom 18.6 + 44.8 1.9 + 1.9 56.3 43.6 58.0 75.4 30.9 47.8 48.3
TinyCLIP-RN30M [68] LAION-400M [51] 15.2B ResNet-30M Custom 29.6 + 54.2 2.6 + 2.6 59.1 45.7 61.5 80.1 33.8 51.6 50.2
TinyCLIP-40M/32 [68] LAION-400M [51] 15.2B ViT-40M/32 Custom 39.7 + 44.5 3.0 + 1.9 59.8 46.5 59.1 76.1 33.5 48.7 51.2
MobileCLIP-S0 DataCompDR-1B 13B MCi0 MCt 11.4 + 42.4 1.5 + 1.6 67.8 55.1 67.7 85.9 40.4 58.7 58.1
OpenAI-RN50 OpenAI-400M [47] 13B ResNet-50 Base 38.3 + 63.4 3.3 + 3.3 59.8 45.1 57.4 80.0 28.5 48.8 48.1
TinyCLIP-61M/32 [68] LAION-400M [51] 15.2B ViT-61M/32 Custom 61.4 + 54.0 4.3 + 2.6 62.4 48.7 62.6 78.7 36.5 52.8 53.0
@@
15.8B ViT-63M/32 Custom (-) (-) 64.5 (-) 66.0 84.9 38.5 56.9 (-)
MobileCLIP-S1 DataCompDR-1B 13B MCi1 Base 21.5 + 63.4 2.5 + 3.3 72.6 60.7 71.0 89.2 44.0 62.2 61.3
OpenAI-RN101 OpenAI-400M [47] 13B ResNet-101 Base 56.3 + 63.4 4.3 + 3.3 62.3 48.5 58.0 79.0 30.7 49.8 50.3
OpenAI-B/32 OpenAI-400M [47] 13B ViT-B/32 Base 86.2 + 63.4 5.9 + 3.3 63.3 48.5 58.8 78.9 30.4 50.1 52.5
LAION-B/32 LAION-2B [52] 32B 65.7 51.9 66.4 84.4 39.1 56.2 54.8
DataComp-B/32 DataComp-1B [18] 13B 69.2 55.2 61.1 79.0 37.1 53.5 58.0
DataComp-B/32-256 DataComp-1B [18] 34B ViT-B/32-256 Base 86.2 + 63.4 6.2 + 3.3 72.8 58.7 64.9 84.8 39.9 57.9 60.9
MobileCLIP-S2 DataCompDR-1B 13B MCi2 Base 35.7 + 63.4 3.6 + 3.3 74.4 63.1 73.4 90.3 45.4 63.4 63.7
VeCLIP-B/16 [32] WIT-200M 6.4B ViT-B/16 Base 86.2 + 63.4 11.5 + 3.3 64.6 (-) 76.3 91.1 48.4 67.2 (-)
OpenAI-B/16 WIT-400M [47] 13B Base 86.2 + 63.4 11.5 + 3.3 68.3 55.9 67.7 85.9 40.4 58.7 58.1
LAION-B/16 LAION-2B [52] 34B Base 86.2 + 63.4 11.5 + 3.3 70.2 56.6 69.8 86.3 42.3 59.4 58.7
EVA02-B/16 Merged-2B [55] 8B Base 86.2 + 63.4 (-) 74.7 59.6 71.5 86.0 42.2 58.7 58.9
DFN-B/16 DFN-2B [16] 13B Base 86.2 + 63.4 11.5 + 3.3 76.2 62.3 69.1 85.4 43.4 60.4 60.9
DataComp-B/16 DataComp-1B [18] 13B Base 86.2 + 63.4 11.5 + 3.3 73.5 60.8 69.8 86.3 42.3 59.4 61.5
SigLIP-B/16 [77] Webli-1B 40B Custom 92.9 + 110.3 9.9 + 5.8 76.0 61.0 74.7 89.1 47.8 65.7 62.3
MobileCLIP-B DataCompDR-1B 13B Base 86.3 + 63.4 10.4 + 3.3 76.8 65.6 77.3 91.4 50.6 68.8 65.2
MobileCLIP-B (LT) DataCompDR-1B 39B Base 86.3 + 63.4 10.4 + 3.3 77.2 66.1 76.9 92.3 50.0 68.7 65.8
Table 7: MobileCLIP family of models has the best average performance at various latencies. Retrieval performances are reported @1. Last column shows average performance on 38 datasets as in OpenCLIP [29]. Models are grouped by their total latency in increasing order and by performance within each group. “Base” refers to standard CLIP Transformer-based [63] text encoder with 12 layers, and “Custom” stands for customized text encoder used in the respective method. For TinyCLIP-63M/32 and EVA02-B/16, we were unable to reliably benchmark models. Note: EVA02-B/16 [55] uses MIM pretrained weights for its vision encoder and OpenCLIP-B pretrained weights for its text encoder. TinyCLIP models use advanced weight initialization methods utilizing OpenCLIP models trained on LAION-2B[52] dataset. All other models, including ours are trained from scratch. “(LT)” refers to longer training schedule, described in detail in Appendix I.
@@ 5.5 Retrieval Performance Analysis
Table 8: Performance on ARO benchmark. All the models use ViT-B/16 as image encoder and the Base text encoder. For VG Rel. and VG Attr. datasets, Macro Acc. is reported and for Flickr30k-Order and COCO-Order recall@1 is reported following [75].
@@ 6 Conclusion
Acknowledgments
We thank Jason Ramapuram, Vaishaal Shankar, and Russ Webb for their valuable feedback and discussions. We also thank Albin Madappally Jose for support in dataset creation and processing. Finally, we would like to thank the entire Machine Learning Research team at Apple for helpful discussions and assistance with the infrastructure.
@@ Acknowledgments
Appendix A Image Encoder Configurations
In our work, we introduce 3 stage configurations for FastViT architecture that substantially improves the model with limited impact on latency. The three configurations are described in Tab. 9. Comparison of our image encoders with FastViT image encoder when trained on ImageNet-1k dataset in a supervised setting (described in Appendix B) is shown in Fig. 7.
Variant $\{C_{1},C_{2},C_{3},C_{4}\}$$\{L_{1},L_{2},L_{3},L_{4}\}$
MCi0 $\{64,128,256,512\}$$\{2,6,10,2\}$
MCi1 $\{64,128,256,512\}$$\{4,12,20,4\}$
MCi2 $\{80,160,320,640\}$$\{4,12,24,4\}$
Table 9: Configurations of MCi.
Appendix B Experimental Setup
Additional details of our training and evaluation are provided in this section. Table 12 summarizes the hyperparameters we used to train MobileCLIP-B on DataCompDR-1B. For other variants of MobileCLIP (S0, S1, and S2) we use the same hyperparameters except using $\lambda=1.0$. For experiments on DataCompDR-12M we use global batch size of 8192. All models trained on DataComp(-DR) use strong image augmentation unless stated otherwise.
For our ensemble distillation ablations in Appendix D, we use 32 total A100 GPUs but we use the same global batch size of 8192 as our other ablations. We also use a smaller uniformly sampled DataComp-8M for ablations in Appendices D and C that results in a slightly lower performance than DataCompDR-12M used for the rest of ablations.
The seen samples reported for DataCompDR is a triplet of one randomly augmented image, one ground-truth caption, and one randomly picked synthetic caption. The reported number of iterations is the number of seen samples divided by the global batch size.
For ImageNet-1k experiments, we follow the training recipe prescribed in [59, 38], i.e. the models are trained for 300 epochs using AdamW optimizer with weight decay of 0.05 and peak learning rate $10^{-3}$ for a total batch size of 1024. The number of warmup epochs is set to 5 and cosine schedule is used to decay the learning rate. The teacher model for distillation is RegNetY-16GF [48] Our implementation uses Timm library [66] and all the models were trained on single machine with 8$\times$×NVIDIA A100 GPUs. The hyperparameters for the three variants of MCi are detailed in Tab. 10. The performance of MCi variants is detailed in Tab. 11 and compared against recent state-of-art efficient architectures. MCi obtains the best trade-off amongst recent efficient architectures as seen in Fig. 7.
@@ Appendix B Experimental Setup
Appendix C Image Augmentation
@@
Appendix D CLIP Ensembles
@@
Appendix E Ablations on Lossy Compressions
In general, the storage size of datasets depends on the file format and the tradeoff between load time and the compression rate. In Tab. 4(c) we presented the storage sizes for DataCompDR-12M and DataCompDR-1B with BFloat16 compression of the embeddings. In this section, we further analyze the storage reduction by i) reducing the number of augmentations, and ii) lossy compression of embeddings.
@@ Appendix E Ablations on Lossy Compressions
Image Text Syn. Aug. Params Text Emb. Image Emb. BFloat16 Sparsity Size (GBs)
@@
Appendix F Hybrid Text Encoder
@@
For the hybrid design, we use depth-wise 2D convolutional layers. We reshape the 3 dimensional input tensor to (BC1S) format, i.e. (Batch Size, Channel Dim., 1, Seq. length) before feeding the tensor to the convolutional layer. For CLIP, the sequence length is set to 77. The depth-wise convolutions enable interactions between tokens across the sequence. The FFN layers enable interactions between token’s channel dimensions. Since the convolution layer is 2D, we simply reuse the reparameterization process described in [62].
@@
Image Enc. Dataset # Image Enc. Params (M) Latency (ms) (img+txt) 0-shot IN-val $\mathbf{\Delta}$
MobileNetv3-L DataComp-12M 4.9 1.1 + 3.3 34.1
DataCompDR-12M (Ours) 44.7 $\uparrow$+10.6
ViT-T/16 DataComp-12M 5.6 3.0 + 3.3 32.9
DataCompDR-12M (Ours) 44.1 $\uparrow$+11.2
ResNet-50 DataComp-12M 24.6 2.6 + 3.3 40.4
DataCompDR-12M (Ours) 51.9 $\uparrow$+11.5
FastViT-MA36 DataComp-12M 43.5 4.3 + 3.3 45.2
DataCompDR-12M (Ours) 58.9 $\uparrow$+13.7
Table 19: DataCompDR-12M vs. DataComp-12M. All the models are trained for 30k iterations ($\sim$ 0.24B seen samples).
Appendix G Performance of other models on DataCompDR-12M
In Tab. 19, we compare performance of CLIP models with different sized image encoders when trained on DataCompDR-12M. All models enjoy significant accuracy improvement when trained on DataCompDR-12M with no training overhead. For example, even the smallest model like MobileNetV3-L with only 4.9M parameters obtains a significant 10.6% improvement in zero-shot IN-val performance.
Appendix H Extended Results
@@
MobileCLIP-B 76.8 58.7 89.6 41.4 64.5 69.8 69.4 77.3 94.4 96.7 91.4 99.1 99.9 50.6 74.9 82.9 68.8 88.3 92.9
MobileCLIP-S2 74.4 49.3 87.0 46.9 62.2 66.8 66.6 73.4 92.3 95.6 90.3 98.9 99.6 45.4 70.1 79.0 63.4 85.1 91.4
MobileCLIP-S1 72.6 40.3 84.7 50.5 60.3 64.9 63.4 71.0 91.3 95.3 89.2 98.0 99.5 44.0 68.9 77.7 62.2 84.3 90.1
MobileCLIP-S0 67.8 26.5 78.6 53.8 55.5 59.9 55.9 67.7 88.8 93.3 85.9 97.1 98.8 40.4 66.0 75.9 58.7 81.1 88.2
@@
VeCLIP-B/16 [32] 64.6 (-) (-) (-) (-) 57.7 (-) 76.3 93.5 96.4 91.1 98.5 99.7 48.4 73.3 81.8 67.2 87.3 92.7
TinyCLIP-63M/32 [68] 64.5 22.8 74.1 (-) 50.8 55.7 31.2 66.0 (-) (-) 84.9 (-) (-) 38.5 (-) (-) 56.9 (-) (-)
CLIPA-B/16 [34] 63.2 26.8 73.2 (-) 48.7 55.6 44.3 58.3 (-) (-) 75.3 (-) (-) 35.2 (-) (-) 53.1 (-) (-)
@@
Appendix I Long training
In Tab. 21 we provide results for training MobileCLIP-B on more than 13B seen samples. We explore continuing the training of MobileCLIP-B to reduce the cost of training from scratch. Recently, [19] has shown that large scale CLIP models can be continually pretrained as the data distribution varies with time. We utilize some of their recipes for continual training where we initialize the training with a model previously trained with cosine or constant learning rate schedule and restart the training on DataCompDR-1B. We utilize a short warmup to stabilize the training and then use another constant or cosine learning rate schedule with maximum and minimum values equal to the original training. We train using 64 nodes with 8xA100-80GB GPUs and a per-GPU batch size of either 128 or 256. One seen sample for DataCompDR is a triplet of one randomly augmented image, one ground-truth caption, and one randomly picked synthetic caption. Number of iterations is the number of seen samples divided by the global batch size. Note that training wall-clock time is the same for DataCompDR vs DataComp (Tab. 4(d)).
Compared with our initial training on 13B seen samples, our long training with 39B total seen samples achieves 0.6% improvement in average performance on 38 datasets as well as 0.4% improvement in zero-shot IN-val accuracy. We reach similar improvements in average performance on 38 datasets (0.4%) with only 18B total seen samples by continuing our original training on 13B seen samples with a short training using Cosine(40k, 131k, 2k).
LR Schedule Seen Samples Zero-shot CLS Flickr30k Ret. COCO Ret. Avg. Perf. on 38
[…]
Cosine(200k, 65k, 2k) 13B 76.8 65.6 77.3 91.4 50.6 68.8 65.2
Const(300k, 65k, 2k) + Cosine(40k, 131k, 2k) 25B 77.1 65.8 77.0 91.8 50.2 68.7 65.2
Const(300k, 65k, 2k) + Cosine(300k, 65k, 2k) 39B 77.2 66.1 76.9 92.3 50.0 68.7 65.8
Const(200k, 65k, 2k) + Cosine(40k, 131k, 2k) 18B 77.1 65.9 77.0 92.8 50.3 69.1 64.6
Cosine(200k, 65k, 2k) + Cosine(40k, 131k, 2k) 18B 76.8 65.6 76.8 92.1 50.4 69.1 65.6
Cosine(100k, 131, 2k) + Cosine(40k, 131k, 2k) 18B 77.0 65.6 77.2 91.3 50.2 69.2 64.2
Table 21: MobileCLIP-B long and continual training. Retrieval performances are reported @1. Last column shows average performance on 38 datasets as in OpenCLIP [29]. The learning rate schedules are specified as Cosine/Const(num-iterations, global batch-size, warmup-iterations). We highlight numbers within 0.2% of the maximum in each column.